# Complexity Growth and the Krylov-Wigner function

### v1

@@ 5 Discussion
In this work, we have introduced and studied the Krylov-Wigner function, i.e., the discrete Wigner function defined with respect to the Krylov basis. We showed that with an appropriate choice of phases, the Krylov basis minimizes the growth of the negativity of the Wigner function for large $D$ quantum systems. We then studied the negativity of the Krylov-Wigner function in Gaussian random matrix theory and observed broadly three phases in its time evolution: spread, ramp and plateau. Much like circuit complexity Brown_2018 ; Balasubramanian:2019wgd ; Balasubramanian:2021mxo , the KW negativity also keeps growing till a time exponential in $\log\,D$, and then saturates at an exponentially large value. We end with some remarks on the outlook and future directions.
@@ 5.2 Other future directions
Refer to caption Refer to caption
Figure 4: Temperature dependence of the negativity of the KW function for $D=251$.
Refer to caption Refer to caption
@@
where $\beta$ is the inverse temperature. For inverse temperature $\beta\sim 1$, the time-dependence of the negativity is not affected very much (see figure 4 for some preliminary results). In particular, it shows the same spread, ramp and plateau behaviour, with a very similar saturation value. However, as the temperature is lowered, the saturation value of the negativity decreases (although it remains proportional to $\sqrt{D}$). The distinction between the spread and the ramp phases also seems to get washed out at smaller temperatures, although this could be an artefact of working at low Hilbert space dimension. It would be interesting to explore the temperature dependence of the negativity in more detail. It would also be interesting to consider other initial states, such as for example the thermofield double state (see figure 5 for some preliminary results in the infinite temperature case). Here, the saturation time seems to be $O(D)$, as opposed to the $O(\sqrt{D})$ scaling of the saturation time for pure initial states. Finally, it would be very interesting to explore these ideas in toy models for holography such as the SYK or the DSSYK model.
@@ Appendix A Early time approximation for the Wigner function
[…]
Figure 6: A comparison of the Wigner negativity computed from the approximate formula (95) and the numerical result for D = 151.
@@
[…]
@@
[…]

### v2

@@ 5 Discussion
In this work, we have introduced and studied the Krylov-Wigner function, i.e., the discrete Wigner function defined with respect to the Krylov basis. We showed that with an appropriate choice of phases, the Krylov basis minimizes the growth of the negativity of the Wigner function for large $D$ quantum systems. We then studied the negativity of the Krylov-Wigner function in Gaussian random matrix theory and observed broadly three phases in its time evolution. Much like circuit complexity Brown_2018; Balasubramanian:2019wgd; Balasubramanian:2021mxo, the KW negativity also keeps growing till a time exponential in $\log\,D$, and then saturates at an exponentially large value. We end with some remarks on the outlook and future directions.
@@ 5.2 Other future directions
Figure 4: Temperature dependence of the negativity of the KW function for $D=101$.
@@
where $\beta$ is the inverse temperature. For inverse temperature $\beta\sim 1$, the time-dependence of the negativity is not affected very much (see figure 4 for some preliminary results). In particular, it shows the same general behavior, with a very similar saturation value. However, as the temperature is lowered, the saturation value of the negativity decreases (although it remains proportional to $\sqrt{D}$). It would be interesting to explore the temperature dependence of the negativity in more detail. It would also be interesting to consider other initial states, such as for example the thermofield double state (see figure 5 for some preliminary results in the infinite temperature case). Finally, it would be very interesting to explore these ideas in toy models for holography such as the SYK or the DSSYK model.
@@ Appendix A Early time approximation for the Wigner function
$\int_{0}^{\pi}d\theta\,|\theta\rangle\langle\theta|=1.$ (90)
Figure 6: A comparison of the Wigner negativity computed from the approximate formula (94) and the numerical result for D = 199.
@@
$\langle k|e^{-itH}|0\rangle\sim\int_{0}^{\pi}d\theta\,\langle k|\theta\rangle\langle\theta|e^{-itH}|0\rangle=\frac{2}{\pi}\int_{0}^{\pi}d\theta\,e^{-2it\cos\theta}\sin[(k+1)\theta]\sin(\theta).$ (91)
@@
$f_{k}(t)$ $=$ $\int_{0}^{\pi}d\theta\,\sin\theta\sin[(k+1)\theta]e^{2it\cos\theta}$ (93)
$=$ $-\frac{1}{2}\int_{0}^{\pi}d\theta\,\left(\cos[(k+2)\theta]-\cos[k\theta]\right)e^{2it\cos\theta}$